OpenStudy (richyw):

Let C be the curve parametrized by \[x=3sint,\;y=4t,\;z=3cost,\;t\geq 0\] a)if you start at the point (0,0,3) and move \(5\pi\) units along C, you end up at a point \((z,y,z)=(a,b,c)\). Find the coordinates of a,b and c. b)Find the curvatuce of C at the point (a,b,c) found in part a

OpenStudy (richyw):

alright before anyone goes crazy, I have the correct answer for a) I'll quickly explain it as I need it for b (the part I don't get)

OpenStudy (richyw):

so we have \[\vec{r}(t)=\langle 3\sin t,4t,3\cos t\rangle\] then \[\vec{v}(t)=\frac{d\vec{r}}{dt}=\langle3\cos t,\,4\,,-3\sin t\rangle\]and \[u=||\vec{v}(t)||=\sqrt{9\cos^2t+16+9\sin^2t}=5\]

OpenStudy (richyw):

so then with arc-length parametrization I also know \[u=\frac{ds}{dt}=5\] so \(ds=5dt\) and \(s=5t\) so \[\vec{r}(s)=\left\langle 3\sin\left({\frac{s}{5}}\right), 4\left(\frac{s}{5}\right),3\cos{\left(\frac {s}{5}\right)} \right\rangle\]

OpenStudy (richyw):

and therefore \[\vec{r}(5\pi)=\left\langle\sin\pi,\,4\pi,\,3\cos\pi\right\rangle=\left\langle 0,\,4\pi,\,-3\right\rangle\]

OpenStudy (anonymous):

Good.

OpenStudy (richyw):

OK so that's where I am. now I know that \[\kappa=\left|\left|\frac{d\vec{T}}{ds}\right|\right| \] so I'm assuming I need to find \(\vec{T}\) which I though was just \[\frac{v(t)}{u(t)}\]

OpenStudy (anonymous):

Yes, that is correct.\[\mathbf{T}=\frac{\mathbf{r'}}{\|\mathbf{r'}\|}\]But... once you calculate that, you must differentiate with respect to s, not t. I presume you have a method. Show me.

OpenStudy (richyw):

well I got \[v(s)=\frac{1}{5}\left\langle 3\sin(s/5),\,4,\, 3\cos(s/5)\right\rangle\]

OpenStudy (richyw):

which the length is also 5 I think

OpenStudy (richyw):

so I get \[\vec{T}=\frac{1}{25} \left\langle 3sin\left(\frac{s}{5}\right),\,4,\,3\cos \left(\frac{s}{5}\right)\right\rangle\]

OpenStudy (richyw):

but this is incorrect.

OpenStudy (richyw):

oops negative for the z component. it's the fraction put front that's wrong though

OpenStudy (anonymous):

I keep getting myself confused. I used the parameter t. Your method is confusing to me, sorry.

OpenStudy (anonymous):

The curve is a helix. The curvature at any point is 3/25

OpenStudy (richyw):

OK I just figured it out I think. thanks a lot!

OpenStudy (richyw):

yup I get the answer you gave now! case closed haha

OpenStudy (anonymous):

Now calculate the torsion!